# Uncertainty in predicting the stochastic gravitational wave background from compact binary coalescences

### v2

@@ Abstract
The stochastic gravitational-wave background from compact binary coalescences is expected to be the first detectable stochastic signal via cross-correlation searches with terrestrial detectors. It encodes the cumulative merger history of stellar-mass binaries across cosmic time, offering a unique probe of the high-redshift Universe. However, predicting the background spectrum is challenging due to numerous modeling choices, each with distinct uncertainties. In this work, we present a comprehensive forecast of the astrophysical gravitational-wave background from binary black holes, binary neutron stars, and neutron star-black hole systems. We systematically assess the impact of uncertainties in population properties, waveform features, and the modeling of the merger rate evolution. By combining all uncertainties, we derive credible bands for the background spectrum, spanning approximately an order of magnitude in the fractional energy density. These results provide thorough predictions to facilitate the interpretation of current upper limits and future detections.
@@ I Introduction
The gravitational-wave observatories LIGO [1], Virgo [2], and KAGRA [3] are gradually revealing the population of merging stellar-mass compact binaries in the Universe [4, 5, 6, 7, 8]. By directly observing hundreds of binary black hole (BBH) mergers, these detectors have provided unprecedented insights into the population of compact objects [9, 10, 11]. While individual detections provide detailed information about the local population, a vast number of mergers remain undetected, either because they occur at high redshifts or because their signals fall below the sensitivity threshold of current instruments. The superposition of these unresolved compact binary coalescences (CBCs) produce a persistent stochastic gravitational-wave background (SGWB) [12, 13], which encodes the integrated history of binary mergers across cosmic time.
@@
The search for a SGWB of astrophysical origin complements the direct observations of individual events, as it provides access to the high-redshift Universe [14, 15, 16, 17, 18, 19]. While this work focuses on the SGWB produced by CBCs, other astrophysical and cosmological sources may also contribute to the background. These include core-collapse supernovae [20, 21, 22], rotating neutron stars [23, 24, 25, 26], primordial black holes [27, 28, 29] and various early-universe phenomena [30, 31, 32, 33, 34]. Forecasts of the astrophysical gravitational-wave background (AGWB) from CBCs have evolved significantly over time. Early predictions, prior to the first gravitational-wave (GW) detections, relied on simple models of binary populations [35, 36, 37, 38, 39, 40, 41, 42]. More recent studies incorporate population synthesis models [43, 44, 45] or observational data into the projections [18, 11, 46, 47, 48]. Despite these advances, predicting the AGWB spectrum remains challenging due to the numerous astrophysical inputs involved, many of which carry substantial uncertainties. These include the mass and spin distribution of the compact binary population, as well as the redshift-dependent merger rate. In particular, modeling the merger rate history is complicated by factors that are difficult to capture with simple prescriptions and vary across different formation channels [49, 50, 51, 45, 52]. Among these are the delay time between binary formation and merger, which is sensitive to the underlying stellar evolution and binary interaction physics, and the metallicity of the progenitor environment, which affects both the formation efficiency and the physical properties of the resulting compact objects [53, 54, 55, 56, 57, 58, 59]. This work focuses on two key aspects of the AGWB from CBCs: (1) accurately modeling the AGWB while properly accounting for the various uncertainties involved in its calculation, and (2) systematically analyzing the impact of individual modeling assumptions on the resulting background spectrum. Our aim is to provide a comprehensive understanding of how different physical inputs and modeling choices shape the AGWB spectrum, and to quantify the credible range within which the background is expected to lie.
@@
We begin by introducing a fiducial model for the AGWB from three classes of CBCs: BBH, BNS, and neutron star–black hole (NSBH) systems. These models reflect the current best estimates of the local binary population, as characterized in Ref. [11], and adopt the merger rate evolution used in the forecast of Ref. [18]. To compute the AGWB spectrum, we evaluate the fractional energy density via Monte Carlo integration over the relevant source parameters, employing state-of-the-art waveform models. We then investigate how the uncertainties in the mass and spin distribution propagates into uncertainties in the AGWB spectrum. We also show how different population models influence the spectral shape and amplitude of the background. Additionally, we examine the impact of waveform features beyond the leading-order quadrupole approximation, including higher-order GW modes in BBH mergers and tidal effects in BNS systems, both of which can significantly alter the spectral features of the background.
@@
Finally, we combine the ensemble of merger rate histories with the uncertainties in the population models to construct credible bands for the AGWB spectrum from BBH, BNS and NSBH. These bands represent the range of plausible background amplitudes and span approximately an order of magnitude in $\Omega_{\mathrm{GW}}(f)$. Compared to most estimates in the literature, our bands are notably broader, this is mainly because of the inclusion of uncertainties in the merger rate modeling. With the upcoming observing runs of the LVK collaboration, cross-correlation searches are expected to start directly constraining the most optimistic predictions for the AGWB.
This paper is structured as follows: In Section II, we introduce the fiducial model of the AGWB from compact binaries, which serves as a reference for evaluating the impact of different modeling choices. Section III discusses uncertainties arising from population models, including modeling choices such as incorporating higher order waveform modes in computing the BBH background and tidal effects for the BNS background. In Section IV, we analyze the individual and combined uncertainties associated with modeling the merger rate history. In Section V, we combine all uncertainties and present the final credible bands for the AGWB from the three CBC classes. Finally, we conclude with a discussion and outlook in Section VI.
@@ II Fiducial Model of the astrophysical background from CBC
A Gaussian, isotropic, unpolarized and stationary GWB is characterized by the fractional energy density spectrum [60, 61]
@@
where $d\rho_{\mathrm{GW}}$ denotes the GW energy density contained in the frequency interval $f+df$ and $\rho_{c}=\frac{3H_{0}^{2}c^{2}}{8\pi G}$ is the critical energy density required today to close the universe, and $H_{0}$ is the Hubble constant. In general, the stochastic GWB from an astrophysical population can be computed using the following expression [12, 37]:
@@ II.1 Fiducial binary populations
For the BBH population, we adopt the best-fit mass distribution inferred from GWTC-3 based on the confident detections with an inverse false alarm rate (iFAR) greater than one year. While our analysis is based on GWTC-3 population inferences, we note that GWTC-4 [62, 63] has since been released, its potential impact on our results is discussed in Section VI. Specifically, we use the preferred power-law plus peak (PLP) model described in the appendix of Ref. [11]. In this model, the primary mass distribution follows a power law with spectral index $a$ between $m_{\mathrm{min}}$ and $m_{\mathrm{max}}$ with an additional Gaussian peak with mean $m_{\mu}$ and width $m_{\sigma}$. The mixing fraction $\lambda_{p}$ determines the relative contribution of the Gaussian peak and the power law component. Additionally, the parameter $\delta_{m}$ governs the smoothing at the lower end of the mass spectrum. The mass ratio $q=m_{2}/m_{1}$ distribution follows a power law with spectral index $\beta_{q}$. The hyperparameters defining the fiducial PLP model are set to: $a=3.5$, $_{\mathrm{min}}=5.1\,M_{\odot}, mm$$_{\mathrm{max}}=88\,M_{\odot}$, $\lambda_{p}=0.038$, $m_{\mu}=34\,M_{\odot}$, $m_{\sigma}=4.6\,M_{\odot}$, $\delta_{m}=5.0\,M_{\odot}$ and $\beta_{q}=1.1$, based on the median values from the population inference presented in Ref. [11]. Given that the observed BBH population exhibits low effective inspiral spins [11, 64], we assume non-spinning BBHs in our fiducial model. Nonetheless, in Section III.2 we explore the impact of alternative spin distributions.
Due to the small number of BNS events compared to BBH events, the BNS mass distribution remains poorly constrained. For our fiducial model, we follow Ref. [18] and assume a uniform distribution for both component masses, $m_{1}$ and $m_{2}$, in the range $1.0\,M_{\odot}$ to $2.5\,M_{\odot}$. Tidal effects due to the matter in neutron stars are neglected for the fiducial model. Additionally, we assume non-spinning BNS systems, in agreement with expectations from Galactic BNS spin measurements [65]. The mass distribution of NSBH is similarly uncertain. Following Ref. [11], we adopt a uniform distribution of neutron star masses between $1\,M_{\odot}$ and $2.5\,M_{\odot}$, and a logarithmically uniform distribution for black hole masses between $5\,M_{\odot}$ and $50\,M_{\odot}$.
@@ II.2 Binary merger rates
where we set the values $\gamma=2.6$, $z_{\mathrm{peak}}=2.2$ and $\kappa=6.2$, according to Madau-Fragos [67]. With redshift the rate first increases with the power-law index $\gamma$, peaks at $z_{\mathrm{peak}}$ and decreases again for high redshifts with negative power-law index $\kappa$. Although Ref. [18] employed a different SFR parametrization following Ref. [68, 69], we demonstrate later that the choice of SFR model has only a minor impact on the resulting GWB.
@@
and normalized to the local merger rate inferred from GW observations [11]. For BBHs modeled with the PLP mass distribution, the rate is best constrained at a redshift $z=0.2$ with a value of $28.3\,\mathrm{Gpc^{-3}\mathrm{yr}^{-1}}$. For the fiducial mass distribution of BNS and NSBH systems, the reported rates at $z=0$ are $105.5\,\mathrm{Gpc^{-3}\mathrm{yr}^{-1}}$ and $32.0\,\mathrm{Gpc^{-3}\mathrm{yr}^{-1}}$, respectively. The time delay distribution $p(t_{d})$ is typically assumed to be log-uniform, with a maximum delay equal to the Hubble time and a minimal delay of a few Myr. In our fiducial model, we adopt a minimum delay of 50 Myr for BBHs and 20 Myr for both BNS and NSBH systems. Additionally, when computing the BBH merger rate, we weight the SFR at the formation redshift $z_{f}$ by the fraction of stellar formation occurring below a metallicity threshold $Z<Z_{\mathrm{th}}$. This accounts for the enhanced efficiency of BBH formation in low-metallicity environments [58, 70]. The metallicity-dependent efficiency is modeled following Ref. [71], as also adopted in Refs. [18, 11, 59]:
@@
where $\hat{\Gamma}$ denotes the incomplete gamma function. For the fiducial BBH model, we assume $Z_{\mathrm{th}}=0.1\,Z_{\odot}$, different thresholds and alternative metallicity models are investigated in later sections.
@@ II.3 CBC background
The energy density spectrum emitted by a single CBC source is given by [12]:
@@
where $h_{+}(f)$ and $h_{\times}(f)$ are the plus and cross polarizations of the gravitational waveform in the frequency domain and $r$ is the distance at which the waveform is evaluated. Note that the distance dependence cancels out, as $h_{+}\,,h_{\times}\sim r^{-1}$. With all ingredients in place, we compute the GWB for the three source classes using Eq. 2. Since this involves computing a complicated high-dimensional integral, we employ a Monte-Carlo technique [72]. For each variable parameter, we draw a large number $N$ of samples from their respective probability distributions and estimate the GWB via:
@@
We assume a flat $\mathrm{\Lambda CDM}$ cosmological model and set the following parameters according to the Planck 2018 results [77]: $H_{0}=67.4\,\mathrm{km}\,\mathrm{Mpc}^{-1}\mathrm{s}^{-1}$, $\Omega_{M}=0.315$ and $\Omega_{\Lambda}=1-\Omega_{M}$. In Fig. 1, we present the fractional energy density spectrum of the GWB from the three classes of CBCs, along with their combined contribution, based on the fiducial models. The dependence of $\Omega_{\mathrm{GW}}(f)$ on cosmological parameters arises primarily through the $1/H_{0}^{2}$ scaling. Additionally, the SFR history is affected via changes in the comoving volume element, while variations in $\Omega_{M}$ have a negligible effect. If instead, we would adopt a cosmology based on the SH0ES measurement of the Hubble constant [78], with $H_{0}=73.04\,\mathrm{km}\,\mathrm{s}^{-1}\mathrm{Mpc}^{-1}$ and $\Omega_{M}=0.262$, this would result in a reduction of the amplitude of $\Omega_{\mathrm{GW}}$ by approximately 20%.
@@ III Investigating the uncertainty in the binary population
We investigate the uncertainty in the GWB energy spectrum arising from current constraints on compact binary mass and spin distributions. For BNS systems, we additionally examine the impact of different equation of states (EOSs) on the resulting energy spectrum.
@@ III.1 BBH mass models
In Section II.1, we introduced the PLP mass distribution model, which provides the best fit to the observed BBH population in GWTC-3 [11]. For our fiducial model we adopted a single realization of this model, using the median values of the hyperparameters inferred from the population analysis. To accurately capture the uncertainties and correlations among these hyperparameters, we incorporate the full set of publicly available posterior samples obtained from the hierarchical Bayesian analysis of BBH mergers in GWTC-3. For each sample, we compute the corresponding GWB energy spectrum. The resulting ensemble of spectra, shown in Fig. 2, reflects the full uncertainty associated with the mass distribution. Additionally, we highlight the 90% credible band arising solely from the uncertainty in the mass distribution. At lower frequencies, the background amplitude is proportional to the average chirp mass of the BBH population raised to the power of 5/3, where the chirp mass is the combination of component masses $\mathcal{M}=(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$. For the PLP mass distribution model, this average is primarily a function of the spectral index $a$ of the power law and $\lambda_{p}$, the fraction of binaries in the Gaussian peak. A steeper power law leads to more low mass binaries, while a smaller $\lambda_{p}$ reduces the contribution from the Gaussian peak near $34~M_{\odot}$. The exact location and width of the Gaussian peak have only a marginal effect. These two parameters also largely control the shape of the background spectrum near its maximum. While a steeper power law shifts the peak toward higher frequencies, a dominant Gaussian peak tends to broaden it toward lower frequencies.
@@
To contextualize the uncertainty in the background spectrum derived from the GWTC-3 PLP model, we compare its 90% credible band with those obtained using different population models and underlying GW data. Specifically, we consider the credible bands computed using the PLP and broken power law (BPL) models from GWTC-2 [7, 10], as well as the simple power law model adopted in GWTC-1 [6, 9]. This comparison, shown in Fig. 3, illustrates how our understanding of the BBH mass distribution has evolved with the increasing number of GW detections. The broader credible intervals in earlier models reflect the larger uncertainty due to fewer BBH observations. Notably, the two parametric models from GWTC-2, PLP and BPL, result in very similar credible bands.

### v3

@@ Abstract
The stochastic gravitational wave background from compact binary coalescences is expected to be the first detectable stochastic signal via cross-correlation searches with terrestrial detectors. It encodes the cumulative merger history of stellar-mass binaries across cosmic time, offering a unique probe of the high-redshift Universe. However, predicting the background spectrum is challenging due to numerous modeling choices, each with distinct uncertainties. In this work, we present a comprehensive forecast of the astrophysical gravitational wave background from binary black holes, binary neutron stars, and neutron star–black hole systems. We systematically assess the impact of uncertainties in population properties, waveform features, and the modeling of the merger rate evolution. By combining all uncertainties, we derive credible bands for the background spectrum, spanning approximately an order of magnitude in the fractional energy density. These results provide thorough predictions to facilitate the interpretation of current upper limits and future detections.
@@ I Introduction
The gravitational wave observatories LIGO [1], Virgo [2], and KAGRA [3] are gradually revealing the population of merging stellar-mass compact binaries in the Universe [4, 5, 6, 7, 8]. By directly observing hundreds of binary black hole (BBH) mergers, these detectors have provided unprecedented insights into the population of compact objects [9, 10, 11]. While individual detections provide detailed information about the local population, a vast number of mergers remain undetected, either because they occur at high redshifts or because their signals fall below the sensitivity threshold of current instruments. The superposition of these unresolved compact binary coalescences (CBCs) produces a persistent stochastic gravitational wave background (SGWB) [12, 13], which encodes the integrated history of binary mergers across cosmic time.
@@
The search for a SGWB of astrophysical origin complements the direct observations of individual events, as it provides access to the high-redshift Universe [14, 15, 16, 17, 18, 19]. While this work focuses on the SGWB produced by CBCs, other astrophysical and cosmological sources may also contribute to the background. These include core-collapse supernovae [20, 21, 22], rotating neutron stars [23, 24, 25, 26], primordial black holes [27, 28, 29], and various early-Universe phenomena [30, 31, 32, 33, 34]. Forecasts of the astrophysical gravitational wave background (AGWB) from CBCs have evolved significantly over time. Early predictions, prior to the first gravitational wave (GW) detections, relied on simple models of binary populations [35, 36, 37, 38, 39, 40, 41, 42]. More recent studies incorporate population synthesis models [43, 44, 45] or observational data into the projections [18, 11, 46, 47, 48]. Despite these advances, predicting the AGWB spectrum remains challenging due to the numerous astrophysical inputs involved, many of which carry substantial uncertainties. These include the mass and spin distribution of the compact binary population, as well as the redshift-dependent merger rate. In particular, modeling the merger rate history is complicated by factors that are difficult to capture with simple prescriptions and vary across different formation channels [49, 50, 51, 45, 52]. Among these are the delay time between binary formation and merger, which is sensitive to the underlying stellar evolution and binary interaction physics, and the metallicity of the progenitor environment, which affects both the formation efficiency and the physical properties of the resulting compact objects [53, 54, 55, 56, 57, 58, 59]. This work focuses on two key aspects of the AGWB from CBCs: (1) accurately modeling the AGWB while properly accounting for the various uncertainties involved in its calculation, and (2) systematically analyzing the impact of individual modeling assumptions on the resulting background spectrum. Our aim is to provide a comprehensive understanding of how different physical inputs and modeling choices shape the AGWB spectrum, and to quantify the credible range within which the background is expected to lie.
@@
We begin by introducing a fiducial model for the AGWB from three classes of CBCs: BBH, binary neutron star (BNS), and neutron star–black hole (NSBH) systems. These models reflect the current best estimates of the local binary population, as characterized in Ref. [11], and adopt the merger rate evolution used in the forecast of Ref. [18]. To compute the AGWB spectrum, we evaluate the fractional energy density via Monte Carlo integration over the relevant source parameters, employing state-of-the-art waveform models. We then investigate how the uncertainties in the mass and spin distribution propagate into uncertainties in the AGWB spectrum. We also show how different population models influence the spectral shape and amplitude of the background. Additionally, we examine the impact of waveform features beyond the leading-order quadrupole approximation, including higher-order GW modes in BBH mergers and tidal effects in BNS systems, both of which can significantly alter the spectral features of the background.
@@
Finally, we combine the ensemble of merger rate histories with the uncertainties in the population models to construct credible bands for the AGWB spectrum from BBH, BNS, and NSBH systems. These bands represent the range of plausible background amplitudes and span approximately an order of magnitude in $\Omega_{\mathrm{GW}}(f)$. Compared to most estimates in the literature, our bands are notably broader, mainly due to the inclusion of uncertainties in the merger-rate modeling. With the upcoming observing runs of the LVK collaboration, cross-correlation searches are expected to start directly constraining the most optimistic predictions for the AGWB.
This paper is structured as follows: In Sec. II, we introduce the fiducial model of the AGWB from compact binaries, which serves as a reference for evaluating the impact of different modeling choices. Sec. III discusses uncertainties arising from population models, including modeling choices such as incorporating higher-order waveform modes in computing the BBH background and tidal effects for the BNS background. In Sec. IV, we analyze the individual and combined uncertainties associated with modeling the merger rate history. In Sec. V, we combine all uncertainties and present the final credible bands for the AGWB from the three CBC classes. Finally, we conclude with a discussion and outlook in Sec. VI.
@@ II Fiducial Model of the astrophysical background from CBC
A Gaussian, isotropic, unpolarized, and stationary GWB is characterized by the fractional energy density spectrum [60, 61],
@@
where $d\rho_{\mathrm{GW}}$ denotes the GW energy density contained in the frequency interval $f+df$, $\rho_{c}=\frac{3H_{0}^{2}c^{2}}{8\pi G}$ is the critical energy density required today to close the Universe, and $H_{0}$ is the Hubble constant. In general, the stochastic GWB from an astrophysical population can be computed using the following expression [12, 37]:
@@ II.1 Fiducial binary populations
For the BBH population, we adopt the best-fit mass distribution inferred from GWTC-3 based on the confident detections with an inverse false alarm rate greater than one year. While our analysis is based on GWTC-3 population inferences, we note that GWTC-4 [62, 63] has since been released, and its potential impact on our results is discussed in Sec. VI. Specifically, we use the preferred power-law plus peak (PLP) model described in the Appendix of Ref. [11]. In this model, the primary mass distribution follows a power law with spectral index $a$ between $m_{\mathrm{min}}$ and $m_{\mathrm{max}}$ with an additional Gaussian peak with mean $m_{\mu}$ and width $m_{\sigma}$. The mixing fraction $\lambda_{p}$ determines the relative contribution of the Gaussian peak and the power-law component. Additionally, the parameter $\delta_{m}$ governs the smoothing at the lower end of the mass spectrum. The mass ratio $q=m_{2}/m_{1}$ distribution follows a power law with spectral index $\beta_{q}$. The hyperparameters defining the fiducial PLP model are set to $a=3.5$, $_{\mathrm{min}}=5.1\,M_{\odot}, mm$$_{\mathrm{max}}=88\,M_{\odot}$, $\lambda_{p}=0.038$, $m_{\mu}=34\,M_{\odot}$, $m_{\sigma}=4.6\,M_{\odot}$, $\delta_{m}=5.0\,M_{\odot}$ and $\beta_{q}=1.1$, based on the median values from the population inference presented in Ref. [11]. Given that the observed BBH population exhibits low effective inspiral spins [11, 64], we assume nonspinning BBHs in our fiducial model. Nonetheless, in Sec. III.2 we explore the impact of alternative spin distributions.
Due to the small number of BNS events compared to BBH events, the BNS mass distribution remains poorly constrained. For our fiducial model, we follow Ref. [18] and assume a uniform distribution for both component masses, $m_{1}$ and $m_{2}$, in the range $1.0\,M_{\odot}$ to $2.5\,M_{\odot}$. Tidal effects due to the matter in neutron stars are neglected for the fiducial model. Additionally, we assume nonspinning BNS systems, in agreement with expectations from Galactic BNS spin measurements [65]. The mass distribution of NSBH is similarly uncertain. Following Ref. [11], we adopt a uniform distribution of neutron star masses between $1\,M_{\odot}$ and $2.5\,M_{\odot}$, and a logarithmically uniform distribution for black hole masses between $5\,M_{\odot}$ and $50\,M_{\odot}$.
@@ II.2 Binary merger rates
where we set the values $\gamma=2.6$, $z_{\mathrm{peak}}=2.2$, and $\kappa=6.2$, according to Madau-Fragos [67]. With redshift the rate first increases with the power-law index $\gamma$, peaks at $z_{\mathrm{peak}}$ and decreases again for high redshifts with negative power-law index $\kappa$. Although Ref. [18] employed a different SFR parametrization following Ref. [68, 69], we demonstrate later that the choice of SFR model has only a minor impact on the resulting GWB.
@@
and normalized to the local merger rate inferred from GW observations [11]. For BBHs modeled with the PLP mass distribution, the rate is best constrained at a redshift $z=0.2$ with a value of $28.3\,\mathrm{Gpc^{-3}\mathrm{yr}^{-1}}$. For the fiducial mass distribution of BNS and NSBH systems, the reported rates at $z=0$ are $105.5\,\mathrm{Gpc^{-3}\mathrm{yr}^{-1}}$ and $32.0\,\mathrm{Gpc^{-3}\mathrm{yr}^{-1}}$, respectively. The time-delay distribution $p(t_{d})$ is typically assumed to be log-uniform, with a maximum delay equal to the Hubble time and a minimal delay of a few Myr. In our fiducial model, we adopt a minimum delay of 50 Myr for BBHs and 20 Myr for both BNS and NSBH systems. Additionally, when computing the BBH merger rate, we weight the SFR at the formation redshift $z_{f}$ by the fraction of stellar formation occurring below a metallicity threshold $Z<Z_{\mathrm{th}}$. This accounts for the enhanced efficiency of BBH formation in low-metallicity environments [58, 70]. The metallicity-dependent efficiency is modeled following Ref. [71], as also adopted in Refs. [18, 11, 59],
@@
where $\hat{\Gamma}$ denotes the incomplete gamma function. For the fiducial BBH model, we assume $Z_{\mathrm{th}}=0.1\,Z_{\odot}$, and different thresholds and alternative metallicity models are investigated in later sections.
@@ II.3 CBC background
The energy density spectrum emitted by a single CBC source is given by [12]
@@
where $h_{+}(f)$ and $h_{\times}(f)$ are the plus and cross polarizations of the gravitational waveform in the frequency domain and $r$ is the distance at which the waveform is evaluated. Note that the distance dependence cancels out, as $h_{+}\,,h_{\times}\sim r^{-1}$. With all ingredients in place, we compute the GWB for the three source classes using Eq. 2. Since this involves computing a complicated high-dimensional integral, we employ a Monte Carlo technique [72]. For each variable parameter, we draw a large number $N$ of samples from their respective probability distributions and estimate the GWB via
@@
We assume a flat $\mathrm{\Lambda CDM}$ cosmological model and set the following parameters according to the Planck 2018 results [77]: $H_{0}=67.4\,\mathrm{km}\,\mathrm{Mpc}^{-1}\mathrm{s}^{-1}$, $\Omega_{M}=0.315$ and $\Omega_{\Lambda}=1-\Omega_{M}$. In Fig. 1, we present the fractional energy density spectrum of the GWB from the three classes of CBCs, along with their combined contribution, based on the fiducial models. The dependence of $\Omega_{\mathrm{GW}}(f)$ on cosmological parameters arises primarily through the $1/H_{0}^{2}$ scaling. Additionally, the SFR history is affected via changes in the comoving volume element, while variations in $\Omega_{M}$ have a negligible effect. If, instead, we were to adopt a cosmology based on the SH0ES measurement of the Hubble constant [78], with $H_{0}=73.04\,\mathrm{km}\,\mathrm{s}^{-1}\mathrm{Mpc}^{-1}$ and $\Omega_{M}=0.262$, this would result in a reduction of the amplitude of $\Omega_{\mathrm{GW}}$ by approximately 20%.
@@ III Investigating the uncertainty in the binary population
We investigate the uncertainty in the GWB energy spectrum arising from current constraints on compact binary mass and spin distributions. For BNS systems, we additionally examine the impact of different equations of states (EOSs) on the resulting energy spectrum.
@@ III.1 BBH mass models
In Sec. II.1, we introduced the PLP mass distribution model, which provides the best fit to the observed BBH population in GWTC-3 [11]. For our fiducial model we adopted a single realization of this model, using the median values of the hyperparameters inferred from the population analysis. To accurately capture the uncertainties and correlations among these hyperparameters, we incorporate the full set of publicly available posterior samples obtained from the hierarchical Bayesian analysis of BBH mergers in GWTC-3. For each sample, we compute the corresponding GWB energy spectrum. The resulting ensemble of spectra, shown in Fig. 2, reflects the full uncertainty associated with the mass distribution. Additionally, we highlight the 90% credible band arising solely from the uncertainty in the mass distribution. At lower frequencies, the background amplitude is proportional to the average chirp mass of the BBH population raised to the power of 5/3, where the chirp mass is the combination of component masses $\mathcal{M}=(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$. For the PLP mass distribution model, this average is primarily a function of the spectral index $a$ of the power law and $\lambda_{p}$, the fraction of binaries in the Gaussian peak. A steeper power law leads to more low-mass binaries, while a smaller $\lambda_{p}$ reduces the contribution from the Gaussian peak near $34~M_{\odot}$. The exact location and width of the Gaussian peak have only a marginal effect. These two parameters also largely control the shape of the background spectrum near its maximum. While a steeper power law shifts the peak toward higher frequencies, a dominant Gaussian peak tends to broaden it toward lower frequencies.
@@
To contextualize the uncertainty in the background spectrum derived from the GWTC-3 PLP model, we compare its 90% credible band with those obtained using different population models and underlying GW data. Specifically, we consider the credible bands computed using the PLP and broken power-law (BPL) models from GWTC-2 [7, 10], as well as the simple power-law model adopted in GWTC-1 [6, 9]. This comparison, shown in Fig. 3, illustrates how our understanding of the BBH mass distribution has evolved with the increasing number of GW detections. The broader credible intervals in earlier models reflect the larger uncertainty due to fewer BBH observations. Notably, the two parametric models from GWTC-2, PLP and BPL, result in very similar credible bands.